OpenStudy (anonymous):

A pyramid has a square base with side s. The height of the pyramid is 2/3 that of its side. What is the expression for the volume of the pyramid?

OpenStudy (anonymous):

v=2s^2 v=2s^3 v=2/3 s^2 v=2/9 s^3 v=1/3 s^3

OpenStudy (anonymous):

im thinking e

OpenStudy (anonymous):

thats right think